you spent $4 on seven pounds of rice and cereal. Rice costs $0.55 per pound and cereal costs $.70 per pound. How many pounds of rice did you buy?
Answer by lwsshak3(11628)

you spent $4 on seven pounds of rice and cereal. Rice costs $0.55 per pound and cereal costs $.70 per pound. How many pounds of rice did you buy?
***
let x=amt of rice bought
7-x=amt of cereal bought
..
.55x+.70(7-x)=4.00
.55x+4.9-.70x=4
.15x=.90
x=6
How many pounds of rice did you buy? 6